OpenStudy (anonymous):

Helppppppppp(:Write an equation of the line, in point-slope form, that passes through the two given points. points: (–17, 8), (3, —2) (y + 8) = –2(x –17) (y – 8) = (x + 17) (y – 8) = –2(x +17) (y + 8) = (x – 17)

OpenStudy (tkhunny):

You should do these almost in your sleep. With the point (-17,8), look first at the x-coordiante. -17 suggests you shoud have x - (-17) = x + 17 Discard 1 and 4 Likewise, you should have y - 8. Too bad. Both 2 and 3 have that. Now calculate the slope and see which one matches.

OpenStudy (anonymous):

(y – 8) = –2(x +17) that what I got

OpenStudy (anonymous):

(y-y1)/(y2-y1)=(x-x1)/(x2-x1) (y-8)/(-2-8)=(x+17)/(3+17) (y-8)/(-10)=(x+17)/(20) y-8=-2(x+17)

OpenStudy (anonymous):

Thats what I got thank you for the help you guys!!(: